Dominique is a manager for a regional bank. He is being relocated several states away to act as a temporary manager while a new branch is interviewing for a permanent manager. He will leave on May 1, 2018, and will be at the new location for less than one year. He sells his personal residence on April 15, 2018, for $123,000 (adjusted basis $95,000). Upon completion of the assignment, he purchases a new residence for $200,000.

What are Dominique’s realized and recognized gain or loss?



What is Dominique’s basis in the new residence?



Assume that Dominique is transferred out of state and sells his new residence for $230,000 two months later (he is single). What are the realized and recognized gains?
Answer #1

a. What are Dominique’s realized and recognized gain or loss?

Selling Price of residence = $123,000

Adjusted basis = $95,000

Realized gain = $123,000 - $95,000 = $28,000

As per Section 121 of IRC, an individual can exclude capital gains up to $250,000. Hence no gain is to be recognized.

b. What is Dominique’s basis in the new residence?

Basis of new residence = Price of new residence = $200,000

c. Assume that Dominique is transferred out of state and sells his new residence for $230,000 two months later (he is single). What are the realized and recognized gains?

As per Section 121 of IRC an individual can exclude capital gains up to $250,000 if he has used the principal residence for at least two of the five years before the sale.

Since Dominique has used it only for 2 months he will be allowed exclusion proportionately.

Two months = 2 * 30 = 60 days ; Two years = 2 * 365 = 730 days

Allowed exclusion = (60 / 730) * 250,000 = $20,548

Gain on Sale = Selling price - Purchase Price

= $230,000 - $200,000 = $30,000

Taxable Gain = Gain on Sale - Allowed exclusion

= $30,000 - $20,548 = $9,452